solve the homogenbosLS Linear system linear system of of the I folloing DER = x + x2 + 3x3 - 2x2 + 2x3. dx = 3X3
Answer #1

A) Given data- Homogeneous linear system of the following differential equations are = x,txat 3x3 dx de = 2x2 + 2x 3 de = 3Xlising eigen values Solving the Linear System of differential equations det (A-AI) = 0 Глоо À l I = 0 1 0 AndI = [1- 3 0 2-3case 1) - x=1 (A-XI) k=0 (A-1I) k = 0.0 - O (A-II) ka.ro 137 0 0 2 E To I 375t equations are Ok, + 2 kq + 3 k₂ = 0 Ok, t kz tCase (ii) - ^2=2 (A-2I) kara I 37k equations are -k, tkat 3 k z=0 - 3 2kg=0 [kg = 0. ea becomes -k, tka=0 a case fin 73 - 3 -- 2141+k, 3k 3 - 0 > - , + 24320 → , - 2 old Id d - B 23, x C + ke +cke +Ansiven